ATTENZIONE: Lunedì 15 Luglio, dalle 13:00 alle 15:00 sarà effettuato un intervento di manutenzione su agenda.infn.it. Durante tale fascia oraria il servizio potrà risultare non raggiungibile.


ATTENTION: On Monday 15 July, from 1:00 pm to 3:00 pm, maintenance will be carried out on agenda.infn.it. During this time slot, the service may be unreachable.

Eco-gas meeting RPC ATLAS+CMS+EPDT+LCHb/Ship

Europe/Rome

Luce presented status of system.

Chamber are now running with ECO2 after Test beam.

Gianluca proposed to do an argon scan to all chambers before restarting the aging.

Stefania presented the results of EPDT chamber during TB 2024 April.

Clear shift of WP and reduction of max efficiency in 2024 vs 2023. The currents seem to be 10% higher for all mixtures, and rates are higher.

Not clear why reduction of max efficiency. Under study, if wp shift is related to current increase.

The increase of rate is not easy to understand if related to different setup in front of chambers , or something else. Position of dosemeter doesn't help.

Giuliana presented an update on resistivity scan analysinzing in spill and out of spill currents, but no clear differences are found.

 

There are minutes attached to this event. Show them.
    • 3:00 PM 3:20 PM
      weekly report 20m
      Speaker: Luca Quaglia (Istituto Nazionale di Fisica Nucleare)
      • April TB successful, we managed to take all requested runs with all mixtures (STD, ECO2 and ECO3), including all single gap runs with RE11 chamber
      • System has been flushing with ECO2 gas mixture since Apr 23 
      • I took two source off scans with all chambers on Wed April 24 (847/848) and the situation seems to be compatible with the last scan taken before TB (731) only slight increase for RE11 BOT (~10 uA from 50 to 60) and EPDT RPC25 (~10 uA from 50 to 60)
      • Following that I resumed a stability scan (still ongoing) with ID 211 -> All chambers included: EPDT @ 9 kV, RE11 and ALICE (new chamber) @ 10.6 kV and SHiP @ 8.8 kV
      • Current increasing in ALICE (not clear why since the chamber is new) and also in RE11 BOT (already observed in past stability scans) as well as other RE11 gaps. Very small increase observed also in KODEL-H BOT (negligible wrt absorbed current)

      Plan

      • Keep stability ongoing until Wednesday (I can take care of it) and perform source off scans
      • Resume regular shifts
    • 3:20 PM 3:40 PM
      status of April Test beam 20m
      Speaker: Luca Quaglia (Istituto Nazionale di Fisica Nucleare)
    • 3:40 PM 4:00 PM
      preliminary results on EPDT chamber during April 2024 TB 20m
      Speakers: Mattia Verzeroli (Université Claude Bernard Lyon I), Dr Stefania Alexandra Juks
    • 4:00 PM 4:20 PM
      update on resistivity measurement 20m
      Speaker: Giuliana Galati (Istituto Nazionale di Fisica Nucleare)
    • 4:20 PM 4:35 PM
      a.o.b. 15m